OpenStudy (anonymous):

a, b, and c are sides of a triangle, with a, b, and c are integers. how many triangles which has the perimeter same with its area ?

OpenStudy (anonymous):

by using trials i have 2 triangles : (6,8,10) and (5,12,13) but i dont know if any else ....

OpenStudy (anonymous):

@mathmath333 , it is more fun :)

ganeshie8 (ganeshie8):

You want to solve \[a+b+c = \dfrac{ab\sin C}{2} \]

ganeshie8 (ganeshie8):

since the only positive integer values that sin(C) can take are \(\{\frac{1}{2}, 1\}\) we get two diophantine equations whose solutions include all the triangles we seek : \[a+b+c = \dfrac{ab}{2}\\~\\a+b+c = \dfrac{ab}{4}\]

ganeshie8 (ganeshie8):

looks like there are exactly 5 solutions http://en.wikipedia.org/wiki/Equable_shape

OpenStudy (anonymous):

thanks, @ganeshie8